OpenStudy (anonymous):

I can not figure out how to factor, I don't know why I am having so much trouble with it. Can someone explain to me how to simplify ((e^x)(2x^1/2X - (e^x)(x^-1/2))/4x especially how to factor the numerator. Thanks so much

OpenStudy (anonymous):

Is this the original? \[\huge{\frac{e^x(2x^{\frac{1}{2}})(-e^x(x^{\frac{-1}{2}}))}{4x}}\]